# Data-Juicer Sandbox: A Feedback-Driven Suite for Multimodal Data-Model Co-development

### v2

@@ D.5 Text-to-Video Generation
For the second task, text-to-video generation, we adopt the advanced DiT-based models, EasyAnimate (Xu et al., 2024b), which originally integrates diverse datasets totaling 1.2M instances from InternVid (Wang et al., 2023b) (606k), Panda-70M (Chen et al., 2024e) (605k), and MSR-VTT (Xu et al., 2016) (6k). The studied baseline model is trained on a subset of 40k instances, employing LoRA (Hu et al., 2021) for efficiency. As a result, the size of $\mathcal{D}$ is 1.2M, and the size of $\mathcal{D}_{sample}$, the single-OP data pools $\mathcal{D}_{i}$ and multi-OP data pools $\mathcal{D}_{\mathcal{S}}$ are all 40k. Model outputs are assessed using VBench (Huang et al., 2024) across 16 metrics on video quality and video-text matchness.
Our investigation covered 21 OPs, including 13 text-only OPs and 10 video-related multimodal OPs. Analogous to the image-to-text generation, we conduct single-OP and multi-OP combination experiments, in Section 4.2 and Section 4.3 respectively. However, given the reduced relevance of data statistics in video-related OPs, our analysis centers on the correlations among the 16 VBench evaluation metrics. These metrics are clustered into three groups, with the best-performing OP selected from each group.
Through OP combination experiments, we pinpoint the most effective set of OPs. In Section 4.4, we then sample 40k instances from the filtered data pool and repeat the training process for up to 10 epochs. For comparative analysis, we adhere to the method outlined in Section 3.2.3, selecting larger data volumes (80k, 120k, …, up to 400k instances) for single-epoch training. To examine the effectiveness of the derived insights for text-to-video data-model co-development, we finally incorporate them into larger-scale scenarios in Section 4.5.
In terms of the model training details, we adopt the advanced DiT-based EasyAnimate (Xu et al., 2024b) model, which integrates diverse datasets totaling 1.2M instances from InternVid (Wang et al., 2023b) (606k), Panda-70M (Chen et al., 2024e) (605k), and MSR-VTT (Xu et al., 2016) (6k). Baseline experiments are executed on a subset of 40k instances, employing LoRA (Hu et al., 2021) for efficiency. During training, we maintain a video resolution of 256x256, sample every other frame, and randomly select sequences of 16 consecutive frames. The training process involves performing a backward pass for the loss of every 8 samples, with single-OP and OP combination experiments trained on a single GPU with a batch size of 8 for 5k steps, amounting to approximately 16 GPU hours per training run. Experiments for duplicating high-quality data, as well as larger-scale training, are conducted with a batch size of 1 across 8 GPUs. The models employ the Adam optimizer for training, with a learning rate set to $2\times 10^{-5}$, weight decay parameter at $3\times 10^{-2}$, and epsilon configured to $10^{-10}$. Each experiment is repeated twice with random seeds of 42 and 45, respectively.
@@ E.1 Summarization of Sandbox Experiments varying Multiple Scales
width = colspec = Q[140]Q[260]Q[280]Q[269], hlines, vlines, Sandbox Scenario & Image-to-Text Generation Text-to-Video Generation Image-Text Pretraining
Main Effectiveness Evidence Optimal recipe derived from small data pools achieves superior model performance in larger data pools, using only half of baseline’s compute cost. Insights from small data pools results in VBench-Top1 model, transferring across data size, model scales, and model architectures Best recipe identified in the model with fewer FLOPS maintains optimal performance with increased model FLOPS and compute resources, showing clear scaling behaviors in power-law form.
@@
(Different FLOPS)
@@
126k to 200k 40k to 147k and 228k 880k to 2,683k
Compute Scale Range (w.r.t Number of Trained Sample) 1 to 8 Epochs 1 to 10 Epochs 4 to 14 Epochs
@@ E.2 Complete Single-Operator Ranking
In Table 6 and Table 7, we present complete numeric results conducted on individual OP experiments (Section 4.2), from which we can discern some more detailed observations.
@@
In image-text pre-training, we can see that the top three performing operations are image_text_similarity_filter, image_text_matching_filter, and image_nsfw_filter. The first two OPs generate statistics based on auxiliary CLIP and BLIP models, enlightening that the modality alignment degree is critical to the studied pair-wise learning task.
@@ E.3 Diversity Analysis
In this subsection, we delve into the diversity of the data in interested data pools. Here, we confine our focus to statistical analysis of words within text data and compute their entropy. We operate under the assumption that the texts provide an accurate description of the images and videos. Consequently, the diversity inferred from the texts also serves as a proxy for the diversity of the associated images and videos.
@@
where $w$ is a word and $\mathcal{P}$ is the distribution of words in a data pool. As we can see, data pools with higher NSFW scores and lower language scores have higher word entropy, suggesting greater diversity within these data pools.
@@ E.4 Correlation Analysis
First, we adopt the most direct approach by examining the Pearson correlation coefficients between the statistics of OPs, as illustrated in Figures 8 and 9. Intuitively, the associations between the statistics of OPs utilized in image-to-text generation appear to be significantly stronger than those in text-to-video generation. For instance, in image-to-text generation, phrase grounding recall shows a strong positive correlation with text perplexity and special character ratio, while exhibiting a strong negative correlation with the alphanumeric ratio, language score, number of text actions, stopword ratio, and text length. In contrast, in text-to-video generation, we observe relationships primarily among the purely textual OPs, while video-related operators are largely orthogonal to others. Therefore, for image-to-text generation, we categorize OPs into three groups based on the correlation of their statistics and select the optimal OP from each category to create combinations. However, this method does not appear appropriate for text-to-video generation due to the sparser correlations observed.
@@
In the second approach, we categorize the evaluation metrics based on their correlations, as illustrated in Figure 10 for image-to-text generation and Figure 11 for text-to-video generation. In image-to-text generation, several evaluation metrics are highly specific and possess unique characteristics, such as those for OCR and coding tasks, resulting in a greater number of categories upon classification. In contrast, VBench’s evaluation metrics can be broadly divided into several classes, including static video metrics (e.g., subject and background consistency), dynamic metrics (e.g., dynamic degree), and video quality indicators (e.g., aesthetic quality and imaging quality).
@@ E.7 Ablation Study Based on T2V-Turbo
The results are shown in Table 3 representing the enhancements we achieved based on T2V-Turbo (Li et al., 2024a). T2V-Turbo applies LoRA (Hu et al., 2021) to VideoCrafter-2.0 (Chen et al., 2024d) and is trained on the WebVid (Bain et al., 2021) dataset, using VideoCrafter-2.0 as a teacher for distillation and incorporating reinforcement learning with rewards for the generated videos. The loss $L$ of T2V-Turbo is defined as:
@@
    Proposed Data Pool (Enhancement from Data View). Utilizing the optimal data pool identified in Section 4.4, which consists of 147k instances with low video NSFW scores and high frame-text similarities, we replace the WebVid data and train for 6 epochs based on the insight from Section 4.4.
@@ E.8 Full Results on VBench leaderboard
Table 12: Detailed FLOPs comparison for representative T2V models. The factor $\alpha$ indicates a
@@
Backed by our proposed methodology and experiments, our Data-Juicer (DJ, with 228k) model refreshes the state-of-the-art on VBench (Huang et al., 2024), surpassing models like Gen-3 (RunwayML, 2024) and Kling (Kuaishou, 2024), as illustrated in Figure 14 and detailed in Table 3.
@@ E.9 Cost Comparison of all Sandbox Experiments
For the text-to-video experiments, Noting that estimating the FLOPS for some models in Table 3 and Table 8 poses several challenges due to the lack of publicly available information. For example, some of their training codes, and training datasets have not been open-sourced, and specific details such as the resolution and frame rate of training videos are unavailable. Thus, we report the estimated FLOPS for the models with sufficient details about the training samples and open-sourced models below. We utilized the inference code of their released open-source models and measured the FLOPS during inference using the calflops  [^3] https://github.com/MrYxJ/calculate-flops.pytorch tool. We then estimated the training FLOPS by applying a factor of $\alpha$ and taking into account the disclosed quantity of training data. From this we can see that our method achieves strong performance while utilizing fewer FLOPS, especially in comparison to the third-place model, VEnhancer, which also originates from VC2, the same upstream model as ours. The second-place Gen-3 and the fourth-place Kling are omitted as they are proprietary models within commercial products.

### v3

@@ D.5 Text-to-Video Generation
For the second task, text-to-video generation, we adopt the advanced DiT-based models, EasyAnimate (Xu et al., 2024b), which originally integrates diverse datasets totaling 1.2M instances from InternVid (Wang et al., 2023) (606k), Panda-70M (Chen et al., 2024e) (605k), and MSR-VTT (Xu et al., 2016) (6k). The studied baseline model is trained on a subset of 40k instances, employing LoRA (Hu et al., 2022) for efficiency. As a result, the size of $\mathcal{D}$ is 1.2M, and the size of $\mathcal{D}_{sample}$, the single-OP data pools $\mathcal{D}_{i}$ and multi-OP data pools $\mathcal{D}_{\mathcal{S}}$ are all 40k. Model outputs are assessed using VBench (Huang et al., 2024) across 16 metrics on video quality and video-text matchness.
Our investigation covered 21 OPs, including 13 text-only OPs and 10 video-related multimodal OPs. Analogous to the image-to-text generation, we conduct single-OP and multi-OP combination experiments, in Sec. 4.2 and Sec. 4.3 respectively. However, given the reduced relevance of data statistics in video-related OPs, our analysis centers on the correlations among the 16 VBench evaluation metrics. These metrics are clustered into three groups, with the best-performing OP selected from each group.
Through OP combination experiments, we pinpoint the most effective set of OPs. In Sec. 4.4, we then sample 40k instances from the filtered data pool and repeat the training process for up to 10 epochs. For comparative analysis, we adhere to the method outlined in Sec. 3.2.3, selecting larger data volumes (80k, 120k, …, up to 400k instances) for single-epoch training. To examine the effectiveness of the derived insights for text-to-video data-model co-development, we finally incorporate them into larger-scale scenarios in Sec. 4.5.
In terms of the model training details, we adopt the advanced DiT-based EasyAnimate (Xu et al., 2024b) model, which integrates diverse datasets totaling 1.2M instances from InternVid (Wang et al., 2023) (606k), Panda-70M (Chen et al., 2024e) (605k), and MSR-VTT (Xu et al., 2016) (6k). Baseline experiments are executed on a subset of 40k instances, employing LoRA (Hu et al., 2022) for efficiency. During training, we maintain a video resolution of 256x256, sample every other frame, and randomly select sequences of 16 consecutive frames. The training process involves performing a backward pass for the loss of every 8 samples, with single-OP and OP combination experiments trained on a single GPU with a batch size of 8 for 5k steps, amounting to approximately 16 GPU hours per training run. Experiments for duplicating high-quality data, as well as larger-scale training, are conducted with a batch size of 1 across 8 GPUs. The models employ the Adam optimizer for training, with a learning rate set to $2\times 10^{-5}$, weight decay parameter at $3\times 10^{-2}$, and epsilon configured to $10^{-10}$. Each experiment is repeated twice with random seeds of 42 and 45, respectively.
@@ E.1 Summarization of Sandbox Experiments varying Multiple Scales
width = colspec = Q[140]Q[260]Q[280]Q[269]Q[270], hlines, vlines, Sandbox Scenario & Image-to-Text Generation Text-to-Video Generation Image-Text Pretraining Image Captioning Fine-tuning
Main Effectiveness Evidence Optimal recipe derived from small data pools achieves superior model performance in larger data pools, using only half of baseline’s compute cost. Insights from small data pools results in VBench-Top1 model, transferring across data size, model scales, and model architectures. Best recipe identified in the model with fewer FLOPS maintains optimal performance with increased model FLOPS and compute resources, showing clear scaling behaviors in power-law form. Optimal recipe identified with InternVL-2.0 model, leads to consistent and steady performance advantages as the model scale increases from 1B to 26B.
@@
(Different FLOPS) InternVL-2 with 1B,
2B, 4B, 26B parameters
@@
126k to 200k 40k to 147k and 228k 880k to 2,683k 24k to 189k
Compute Scale Range (w.r.t Number of Trained Sample) 1 to 8 Epochs 1 to 10 Epochs 4 to 14 Epochs 1 epoch in all experiments
@@ E.2 Complete Single-Operator Ranking
Table 8: The complete OP ranking, including their statistical dimensions and the improvements relative to the baseline for the image captioning task. We consider three splits with low, middle, and high statistical values for each OP. The baseline used is based on random sampling with equal data volume.
OP-Generated Statistics Average Performance Changes (%)
Data Pool (Low) Data Pool (Mid) Data Pool (High)
Text Length 0.760 -3.125 -11.364
Image Watermark Score -0.637 -0.132 0.477
Character Repetition Ratio 0.449 -0.461 -0.634
Text Perplexity -1.505 0.331 -3.220
Word Repetition Ratio 0.300 -3.125 -11.349
Image Width 0.265 -0.230 -0.112
CLIP Image-Text Similarity -0.920 0.225 -2.101
Phrase Grounding Recall -1.318 0.166 0.188
Image Aesthetics Score -0.300 0.186 -0.243
Image Aspect Ratio 0.171 -0.490 -0.118
Image NSFW Score -0.030 0.062 0.113
Image Height -0.080 -0.030 0.088
Stopword Ratio -3.795 0.084 -4.930
Image Size -0.112 -0.071 0.028
Flagged Word Ratio -0.012 -0.663 -0.429
Language Score -0.154 -1.534 -7.823
Token Number -0.610 -0.987 -9.017
Alphanumeric Ratio -0.795 -4.300 -11.592
Text Action Number -2.633 -0.931 -8.565
Word Number -1.125 -2.078 -9.333
Special Character Ratio -11.194 -4.782 -1.145
Entity Dependency Number -1.443 -1.660 -7.072
BLIP Image-Text Similarity -2.254 -1.446 -6.221
In Table 6, Table 7, and Table 8, we present complete numeric results conducted on individual OP experiments (Sec. 4.2), from which we can discern some more detailed observations.
@@
In image-text pre-training, we can see that the top three performing operations are the ones generating stats of CLIP Image-Text Similarity, BLIP Image-Text Similarity, and Image NSFW Score. The first two OPs generate statistics based on auxiliary CLIP and BLIP models, enlightening that the modality alignment degree is critical to the studied pair-wise learning task.
In image captioning fine-tuning, models trained on shorter captions (low range of Text Length) achieve the best performance. Captions with less repetition on both character and word levels help to train models with better performance. It suggests that in the image captioning task, the quality of captions is the key to better performance.
@@ E.3 Diversity Analysis
In this subsection, we delve into the diversity of the data in interested data pools. From the perspective of quantitative indicators, we confine our focus to statistical analysis of words within text data and compute their entropy. We operate under the assumption that the texts provide an accurate description of the images and videos. Consequently, the diversity inferred from the texts also serves as a proxy for the diversity of the associated images and videos.
@@
where $w$ is a word and $\mathcal{P}$ is the distribution of words in a data pool. As we can see, data pools with higher NSFW scores and lower Language Score have higher word entropy, suggesting greater diversity within these data pools.
We also introduce a more direct way to show the data diversity to users. We use a tagging model to extract core objects, concepts, and domains from the contents of the datasets, including both texts and multimodal content. Then we analyze the domain-specific distribution via word clouds of these tags. We show an example of analyzed word clouds from videos on different data pools of Video NSFW Score in Fig. LABEL:fig:internvl. As we can see, although the data pool with high video NSFW scores is the most diverse one, our model achieves the best performance on the data pool with low video NSFW scores, which is consistent with the argument presented in Sec. 4.2.
@@ E.4 Correlation Analysis
Figure 14: Pearson correlation coefficients for OP statistics in image-text pre-training.
Figure 15: Pearson correlation coefficients for OP statistics in image-captioning oriented fine-tuning.
First, we adopt the most direct approach by examining the Pearson correlation coefficients between the statistics of OPs, as illustrated in Figures 12, 13, 14, and 15. Intuitively, the associations between the statistics of OPs utilized in image-to-text generation appear to be significantly stronger than those in text-to-video generation. For instance, in image-to-text generation, phrase grounding recall shows a strong positive correlation with text perplexity and special character ratio, while exhibiting a strong negative correlation with the alphanumeric ratio, language score, number of text actions, stopword ratio, and text length. In contrast, in text-to-video generation, we observe relationships primarily among the purely textual OPs, while video-related operators are largely orthogonal to others. Therefore, for image-to-text generation, we categorize OPs into three groups based on the correlation of their statistics and select the optimal OP from each category to create combinations. However, this method does not appear appropriate for text-to-video generation due to the sparser correlations observed.
@@
In the second approach, we categorize the evaluation metrics based on their correlations, as illustrated in Fig. 16 for image-to-text generation and Fig. 17 for text-to-video generation. In image-to-text generation, several evaluation metrics are highly specific and possess unique characteristics, such as those for OCR and coding tasks, resulting in a greater number of categories upon classification. In contrast, VBench’s evaluation metrics can be broadly divided into several classes, including static video metrics (e.g., subject and background consistency), dynamic metrics (e.g., dynamic degree), and video quality indicators (e.g., aesthetic quality and imaging quality).
@@ E.7 Ablation Study Based on T2V-Turbo
The results are shown in Table 3 representing the enhancements we achieved based on T2V-Turbo (Li et al., 2024a). T2V-Turbo applies LoRA (Hu et al., 2022) to VideoCrafter-2.0 (Chen et al., 2024d) and is trained on the WebVid (Bain et al., 2021) dataset, using VideoCrafter-2.0 as a teacher for distillation and incorporating reinforcement learning with rewards for the generated videos. The loss $L$ of T2V-Turbo is defined as:
@@
    Proposed Data Pool (Enhancement from Data View). Utilizing the optimal data pool identified in Sec. 4.4, which consists of 147k instances with low video NSFW scores and high frame-text similarities, we replace the WebVid data and train for 6 epochs based on the insight from Sec. 4.4.
@@ E.8 Full Results on VBench leaderboard
Table 13: Detailed FLOPs comparison for representative T2V models. The factor $\alpha$ indicates a constant taking into account the disclosed quantity of training data.
@@
Backed by our proposed methodology and experiments, our Data-Juicer (DJ, with 228k) model refreshes the state-of-the-art on VBench (Huang et al., 2024), surpassing models like Gen-3 (RunwayML, 2024) and Kling (Kuaishou, 2024), as illustrated in Fig. 20 and detailed in Table 3.
@@ E.9 Cost Comparison of all Sandbox Experiments
For the text-to-video experiments, Noting that estimating the FLOPS for some models in Table 3 and Table 12 poses several challenges due to the lack of publicly available information. For example, some of their training codes, and training datasets have not been open-sourced, and specific details such as the resolution and frame rate of training videos are unavailable. Thus, we report the estimated FLOPS for the models with sufficient details about the training samples and open-sourced models below. We utilized the inference code of their released open-source models and measured the FLOPS during inference using the calflops  [^3] https://github.com/MrYxJ/calculate-flops.pytorch tool. We then estimated the training FLOPS by applying a factor of $\alpha$ and taking into account the disclosed quantity of training data. From this, we can see that our method achieves strong performance while utilizing fewer FLOPS, especially in comparison to the third-place model, VEnhancer, which also originates from VC2, the same upstream model as ours. The second-place Gen-3 and the fourth-place Kling are omitted as they are proprietary models within commercial products.
E.10 Iterative Workflows in Sandbox
Data-Juicer Sandbox allows iterative workflows, where data-model co-development is activated iteratively to enable data and model improvements for each other. We conduct a further iterative experiment based on the basic single-operator experiments for the image captioning fine-tuning task in Sec. 4.2, where the full experimental results are shown in Table 8.
Experiment Settings. We regard the single-operator recipes in the basic experiments as the $recipe_{1}$, and their trained models as $checkpoint_{1}$. Based on these first-iteration results, we select the best checkpoint from the best single operator Text Length to conduct a second iteration of single-operator continuous training for all single-operator recipes $recipe_{2}$ and obtain the second-iteration trained models $checkpoint_{2}$. Finally, we evaluate the $checkpoint_{2}$ to find out the impacts of the second-iteration data-model co-develpment on the $checkpoint_{1}$.
Table 14: Top-10 OP performance ranking with 2 iterations, including their statistical dimensions and the improvements relative to the baselines for the image captioning task. The 2nd iteration is continuously trained on the top-1 OP (Text Length) checkpoint in the 1st iteration and the improvements are relative to it as well. The OPs in bold in the 2nd iteration also appear in top-10 OPs of the 1st iteration.
{tblr}
cell11 = c=2, cell13 = c=2, vline2 = 1, vline3 = 2-12, hline1-3,13 = -, 1st iter. & 2nd iter.
OP-Generated Statistics Average Performance
Changes (%) OP-Generated Statistics Average Performance
Changes (%)
Text Length 0.760 Flagged Word Ratio 0.607
Image Watermark Score 0.477 CLIP Image-Text Similarity 0.599
Character Repetition Ratio 0.449 Text Perplexity 0.521
Text Perplexity 0.331 Character Repetition Ratio 0.434
Word Repetition Ratio 0.300 Image Aesthetics Score 0.420
Image Width 0.265 Image Size 0.376
CLIP Image-Text Similarity 0.225 Phrase Grounding Recall 0.358
Phrase Grounding Recall 0.188 Image Width 0.320
Image Aesthetics Score 0.186 Image Height 0.285
Image Aspect Ratio 0.171 Image Watermark Score 0.264
Results. As the Table 14 shows, the second iteration continues to improve the model performance. Notably, while the ranking of OP effects changes slightly between iterations, most effective OPs remain consistent: 8 out of the top-10 OPs in the second iteration overlap with those from the first iteration; 7 of top-10 OPs in the first iteration increase their rankings by an average of 6.14 positions and 3 of them decrease by an average of 3.33 positions. Using the same OPs, data pool performance changes, providing actionable insights for dynamic environments.
This also highlights promising future directions and underscores the distinction between our work and DataComp. While DataComp primarily focuses on CLIP pretraining and scaling compute with duplicated data, our data recipes are validated across a broader range of tasks, including text-to-video generation (EasyAnimate, T2V-Turbo), image-to-text pretraining (MGM), post-training (InternVL-2.0) and the iterative co-development.
E.11 Operators beyond Filters
There are diverse types of OPs in Data-Juicer in addition to Filters. In this section, we conduct several experiments on two representative Mappers using MGM-2B on the image-to-text task to verify the effectiveness of operators beyond Filters of the sandbox.
Experiment Settings. We use the same experiment setting of single-operator experiments for the image-to-text task in Sec. 4.2, except that we replace the single-operator recipes of Filters with two new recipes of two Mappers: image_diffusion_mapper: it regenerates a new image for each sample based on the given caption; image_captioning_mapper: it recaptions the existing image in each sample.
Table 15: The performance improvements of two mapper OP, image_diffusion_mapper and image_captioning_mapper, relative to the baseline, which regenerates a new image/caption for each sample using diffusion/BLIP models respectively. The baseline is based on the original raw images.
Mapper OP Average Performance Changes (%)
image_diffusion_mapper 13.45 $\pm$ 11.33
image_captioning_mapper -1.07 $\pm$ 0.65
Results. As the Table 15 shows, images generated by image_diffusion_mapper significantly improve the model performance compared to the original images. In Fig. 21, it highlights a visual example before and after processing with these two Mappers, with the core objects in the captions marked in red. We can observe that diffusion models used in the image_diffusion_mapper effectively locate and better present the key objects (e.g., “setting sun”) in the newly generated image, and it also removes redundant information (e.g., watermarks, link texts). We think this kind of “denoising” capability of image_diffusion_mapper contributes to performance gains. Experiments on other kinds of OPs are left to future work.
Figure 21: An example of image and caption pair before and after being transformed by image_diffusion_mapper and image_captioning_mapper.
E.12 Potential Model Development: Automated Prompt Adjustments
Most parts of this paper focus on the data improvements in data-model co-development. In this section, we present several experiments for automated prompt adjustment to show the potential model development part of the Data-Juicer Sandbox.
Experiment Settings. We select the image captioning fine-tuning task, extend a new model inference hook with about 40 lines of code changes, and update its prompt configuration file. Based on the iterative experiment results in Sec. E.10, we use the top-1 recipes found in two iterations. Then we select 10 different prompts generated by Qwen2.5-max and study the effects of them on trained models in different iterations.
Table 16: The experimental results of automated prompt adjustment on the top-1 recipes found in 2 iterations with 10 different prompts for the image captioning task. The baseline prompt is the default prompt of InternVL2.5. The values that are in bold or underlined are the best or the 2nd best ones in each column.
{tblr}
row2 = c, column4 = c, cell11 = r=2, cell12 = c=2c, cell14 = r=2, cell32 = c, cell33 = c, cell42 = c, cell43 = c, cell52 = c, cell53 = c, cell62 = c, cell63 = c, cell72 = c, cell73 = c, cell82 = c, cell83 = c, cell92 = c, cell93 = c, cell102 = c, cell103 = c, cell112 = c, cell113 = c, cell122 = c, cell123 = c, cell132 = c, cell133 = c, hline1,3-4,14 = -, hline2 = 2-3, Prompt & Average Performance Changes (%) 1st vs. 2nd
1st iter. 2nd iter.
(Baseline) Provide a one-sentence caption for the provided image. 0.760 0.607 >
Provide a brief yet comprehensive caption that highlights the
primary elements and overall mood of the image. 0.944 -0.705 >
Describe the image in one sentence, focusing on the main subject
and its surrounding environment. 0.832 -0.241 >
Summarize the scene depicted in the image with a concise and
descriptive sentence. 0.734 0.148 >
Summarize the content of the image in one clear and accurate
sentence without omitting important details. 0.542 0.186 >
Describe the most important details of the image in one sentence. 0.400 0.130 >
Craft a concise image caption highlighting the main activity
alongside coexisting static elements in the image. 0.083 -1.622 >
What is the main subject of the image, and how would you describe
it in one sentence? -0.026 -0.463 >
If there is movement or activity in the image, summarize it in a
single descriptive sentence. -0.594 0.433 <
Please provide a one-sentence caption that describes the primary
subject’s action and its relationship with the surrounding environment. -1.578 -1.510 <
What stands out most in the image? Capture it in a concise and
striking one-sentence caption. -1.673 -1.300 <
Results. As the Table 16 summarizes, the ranking and performance of these prompts vary significantly across iterations, ranging from -1.673 to +0.944. Some of them beat the baseline prompt, and some of them can enhance or degrade performance in both iterations. It seems that the most suitable prompt is changing during the iterative data-model co-development, which indicates the potential for auto-optimizing model configurations in the context of the co-development of the Data-Juicer Sandbox.